OpenStudy (anonymous):

The equation shows one mole of ethanol fuel being burned in oxygen. Convert the energy released into its equivalent mass. C2 H5 OH (l) + 3 O2 (g) --> 2 CO2 (g) + 3 H2O (l) deltaH = -1418 kJ/mol I am not sure where to begin with this one.

OpenStudy (anonymous):

E=mc^2 ?

OpenStudy (anonymous):

isnt that the only energy-mass relation?

OpenStudy (anonymous):

but first, convert molar mass to kg and then perform the steps

OpenStudy (anonymous):

Should I insert -1.418 * 10^6 for E?

OpenStudy (anonymous):

no. use gram eq. wt. of ethanol to convert deltaH (in per mole) to per kG

OpenStudy (anonymous):

could you show me perhaps?

OpenStudy (anonymous):

1mole C2H5OH = 2(12)+5(1)+16+1 = 46 deltaH = -1418/46 10^6 J/kg \[m = E/c^2\\ m=\frac{-1416\times10^6}{46}\times{1\over (3\times10^8)^2} \]

OpenStudy (anonymous):

but the question asks for energy released by 1 mole . so, \[m=\frac{-1416\times10^3}{(3\times10^8)^2} \]

OpenStudy (anonymous):

-ve sign can be ignored. indicates loss of mass